You are in the process of finding out the value of H&K Company Ltd. The stock currently sells in the market at Tk.52 per share. The dividends paid by the stock over the last five years are:Dividend per share (Tk.) 2020- 3.30 2019- 3.08 2018 -2.86 2017- 2.70 2016-2.50 Looking at these dividend figures you conclude that the pattern of dividend payments has been steady and this pattern will continue in to unforeseeable future. The stock’s year-end market prices from 2016 through 2021 are given below: Year Market price per share (Tk.) 2020 37 2019 34 2018 31 2017 28 2016 26 During the same period the returns on the market as given by the market capitalization rate were as follows. Year Return on market 2019-20 26.50% 2018-19 20.00% 2017-18 19.89% 2016-17 19.74%The risk-free rate of return that is applicable to your investment is the interest rate on three-year term deposit, which is currently 8.5 percent. d. Calculate the required return on H & K stock e. Calculate the value of the stock


1
Expert's answer
2021-05-03T07:04:06-0400

required return on H & K stock


total stock return"= \frac{(P_i-P_o)+D}{P_o}"


where Po = initial stock price

Pi = ending stock price

D = Dividends


P0 =26

Pi=52

D= 14.44


total stock return= "\frac{(52-26)+14.44}{26}"


= 1.555384615

"\frac{1.555384615}{26} \times 100 = 5.9822485%"


stock value = "\frac{DPSI}{r-g}"


where DPSI = expected dividends one year from the present

r= The discount rate or required rate of return on the investment

g=The annual growth rate of dividends in perpetuity


"=\frac{3.52}{0.085-0.0714285713} = 259.3684186" (Tk.)